Consider the parabola given by the equation y = −x^2 + 8x + 3

(a) Find the vertex (x, y)

(b) Solve −x^2 + 8x + 3 = 0. Round your answers to two decimal places.
Answer #1

Equation Y= -x^2 +8x+3

X-axis point = -b/2a = here a=-1 & b= 8

=-8/-2 =4

Y= -(4)^2+8(4)+3= 19

:-vertex (4,19)

(B) equation -x^2 +8x +3 =0

Formula of roots = -b+√D/2a. & -b-√D/2a :- (D= b^2 - 4ac)

=-8+2√19/-2 & -8-2√19/-2

=4-√19 & 4+√19

= -0.35 & 8.35 are the answers